### v1

# DreamDistribution: Prompt Distribution Learning for Text-to-Image Diffusion Models

DreamDistribution: Prompt Distribution Learning for
Text-to-Image Diffusion Models
[…]
${}^{1}$niversity of Southern California ${}^{2}$arvard University
${}^{3}$icrosoft Research Asia ${}^{4}$icrosoft Research Redmond
$*~{}^{*}$=equal contribution as second author ${}^{\dagger}$=equal contribution
[…]
cvprbluergb0.21,0.49,0.74
1 Acknowledgment
This work was supported by the National Science Foundation (award 2318101), C-BRIC (one of six centers in JUMP, a Semiconductor Research Corporation (SRC) program sponsored by DARPA), the Army Research Office (W911NF2020053), and Amazon ML Fellowship. The authors affirm that the views expressed herein are solely their own, and do not represent the views of the United States government or any agency thereof.
Supplementary Material
2 More Implementation Details
In all experiments, we use Stable Diffusion 2.1, which is the latest version. We use the default parameters, including 7.5 guidance scale and 50 denoising steps. For all visual results, we generate images in $768\times 768$ resolution, and for synthetic dataset experiments, we generate $256\times 256$ images to save time and resources. We provide a pseudocode for learning a prompt distribution in Algorithm 1.
Algorithm 1 Training prompt distribution
1:Set of reference images =$\mathcal{I}=\{\mathbf{x}_{i}\}_{i=1}^{N}$
2:Set of learnable prompts $\mathcal{P}^{K}=\{\mathcal{P}_{i}\}_{i=1}^{K}$
3:Text encoder $\mathcal{E}$, noise predictor $\boldsymbol{\epsilon}_{\theta}$, hyperparameter $\lambda$
4:Random initialize all learnable embeddings $\mathbf{V}$ in $\mathcal{P}^{K}$
5:for image $\mathbf{x}\in\mathcal{I}$ do
6:       Sample time step $t$
7:       Sample noise $\boldsymbol{\epsilon}\sim\mathcal{N}(0,\boldsymbol{I})$
8:       $\mathbf{x}_{t}\leftarrow\mathbf{x}$ with noise added based on $\boldsymbol{\epsilon}$ and $t$
9:       Compute $\boldsymbol{\mu_{c}}$ of $\mathcal{E}(\mathcal{P}^{K}$)
10:       Compute $\boldsymbol{\sigma_{c}}$ of $\mathcal{E}(\mathcal{P}^{K})$
11:       for $s\in[S]$ do
12:             Sample $\boldsymbol{\omega}_{s}\sim\mathcal{N}(0,\mathbf{I})$
13:             $\mathcal{L}_{s}\left(\mathcal{P}^{K}\right)=\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\theta}\left(\mathbf{x}_{t},\boldsymbol{\mu_{c}}+\boldsymbol{\omega}_{s}\boldsymbol{\sigma_{c}},t\right)\right\|^{2}_{2}$
14:       end for
15:       $\mathcal{L}\left(\mathcal{P}^{K}\right)=\frac{1}{S}\sum_{s=1}^{S}\mathcal{L}_{s}(\mathcal{P}^{K})$
16:       $\mathcal{L}_{ortho}=\frac{1}{K(K-1)}\sum_{i=1}^{K}\sum_{j=i+1}^{K}\left|\langle\mathcal{E}(\mathcal{P}_{i}),\mathcal{E}(\mathcal{P}_{j})\rangle\right|$
17:       $\mathcal{L}=\mathcal{L}\left(\mathcal{P}^{K}\right)+\lambda\mathcal{L}_{ortho}$
18:       Update learnable embeddings $\mathbf{V}$ in $\mathcal{P}^{K}$ based on $\mathcal{L}$
19:end for
Figure 1: In general, with more prompts, the performance increases in terms of both quality and diversity.
3 Evaluation Set
We show samples of reference images from our evaluation set in Figure 2. Each row shows 8 samples reference images from the same set.
Figure 2: Samples of reference images from our evaluation set
4 Additional Result
4.1 Diverse Image Instance Generation
We show additional image generation results of our method in Figure 3 using the reference images from our evaluation set. Each row is generated using reference images of the corresponding row in Figure 2.
Figure 3: Samples of generated image results using reference images from the evaluation set. Each row is generated using reference images of the corresponding row in Figure 2.
4.2 Text-guided Editing
We show more results on the ability of our method to generate diverse images with text-guided editing in Figure 4 using different sets of reference images.
Figure 4: More results on text-editability of our methods. Left column shows samples of reference images, right columns are generated results with corresponding prompts.
4.3 Scaling Variance for Diversity Control
We show more results on generating images from prompt distribution with scaled standard deviations in Figure 5
Figure 5: More results on scaling standard deviation of learned prompt distribution.
4.4 Composition of Distribution
In Figure 6 we show more results on composition of two different learned prompt distributions with various weights.
Figure 6: More results on composition of multiple prompt distributions using different weights.
5 Ablation study
We additionally ablate $K$, the number of prompts in personalized generation. We randomly select 4 sets of reference images from our evaluation set and compute the average performance based on automatic quality and diversity metrics introduced in main Section 4.1. In Figure 1, we show the effect of $K$ in terms of both generation quality and diversity. We observe a positive correlation between the performance (in terms of both quality and diversity) and the number of prompts. More prompts offer more flexibility for our methods to model a better distribution of prompts, thus enabling the model to encapsulate content better (quality) and adapt to various nuances of the training images (diversity).
6 Synthetic dataset
Figure 7: Column 1 & 2: Top-1 and top-5 accuracy on ImageNet validation set versus using different number of training images to train prompt distribution. Column 3 & 4; Top-1 and top-5 accuracy on ImageNet validation set versus percentage of synthetic images added to the real training set.
6.1 More Analysis on Synthetic Dataset
Number of training images
We experiment with different number of training images. We use randomly selected 10, 100, 500 images per class, as well as all ImageNet training images to train our learnable prompts and generate same size synthetic dataset. From results shown in Figure 7 column 1 & 2, we found that using about 100-500 images per class (7%-38% of the real training set) would be enough to reach high classification accuracy on real validation set, while using more data would not further improve accuracy. For validation sets of ImageNet variants, less training data would obtain higher accuracy due to the domain gap between the real training set and different validation sets.
Mixing synthetic data with real training data
We also experiment with mixing different sizes of synthetic image data with the real training images. We mix additional 20%, 40%, 60%, 80% and 100% synthetic data with real training data, where 100% means the size of the mixed dataset is twice of the size of the ImageNet training set, and the ratio of the number of real images to the number of synthetic images is approximately 1:1. As shown in Figure 7 column 3 & 4, adding more synthetic data would improve the accuracy on ImageNet validation set. On the validation sets of different domains, however, adding more synthetic data would not show significant improving trends on accuracy.
6.2 Implementation detail
For training on ImageNet dataset, we use the same training hyperparameters except for reducing the number of learnable prompts to 10 per class. We train for 5 epochs for training prompt distribution and 300 epochs for training ResNet-50 using generated or mixed dataset. All results are averaged over 3 runs of training using the generated or mixed dataset. For ImageNet-R and ImageNet-A, we only evaluate on the overlapping classes with ImageNet-100.
6.3 Visual result
We show some generated training images using our method and compare them with generated images using solely class names as text prompts for generation in Figure 8. Compared to the images generated using class names, our generated images shows more diversity with different real-world contexts.
Figure 8: Comparison between training images generated using our method and image generated using solely class names as text prompts. For each group of two rows, the top row shows samples of our generated training images, and the bottom row shows the training images generated using class names as text prompts.