### v2

# DreamDistribution: Learning Prompt Distribution for Diverse In-distribution Generation

DreamDistribution: Learning Prompt Distribution for Diverse In-distribution Generation
Brian Nlong Zhao1,  Yuhang Xiao1∗,  Jiashu Xu2∗,  Xinyang Jiang3,  Yifan Yang3,
 Dongsheng Li3,  Laurent Itti1,  Vibhav Vineet4†,  Yunhao Ge1†
1University of Southern California ${}^{2}$arvard University
${}^{3}$icrosoft Research Asia ${}^{4}$icrosoft Research Redmond
${}^{*}$qual Contribution ${}^{\dagger}$qual Advising
1 Introduction
Dreams have long been a source of inspiration and novel insights for many individuals (Edwards et al., 2013; Von Grunebaum & Caillois, 2023). These mysterious subconscious experiences often reflect our daily work and life (Freud, 1921). However, these reflections are not mere replicas; they often recombine elements of our reality in innovative ways, leading to fresh perspectives and ideas. We aim to emulate this fascinating mechanism in the realm of text-to-image generation.
Text-to-image (T2I) generation has recently been popularized due to the astonishing performance of state-of-the-art diffusion models such as Stable Diffusion (Rombach et al., 2021) and DALL·E 2 (Ramesh et al., 2022). Variations of the T2I models have enabled several fascinating applications that allow user to control the generation, such as conditioned generation based on other input modalities (Zhang et al., 2023; Li et al., 2023; Yang et al., 2023), inpainting (Lugmayr et al., 2022; Xie et al., 2023), image editing (Mokady et al., 2023; Brooks et al., 2023). One such interesting application is personalization of T2I models, where user provides some reference images of the same instance (e.g. their pet dog), and the personalized model can generate images based on the references, with the flexibility of text-guided editing for new context. This is generally achieved by associating a token with the personalized concept through fine-tuning the model parameters (Ruiz et al., 2022; Kumari et al., 2023) or newly added learnable token embeddings (Gal et al., 2022; Voynov et al., 2023).
In many cases, however, user may want to personalize T2I generation over a more abstract visual attribute instead of a specific instance-level personalization. For example, a designer may seek inspiration by generating a variety of novel cartoon characters or scenery images following similar visual attributes presented in their previous works. In this case, trying over text prompts is not scalable and hard to get desired result that follows the desired visual attributes. On the other hand, using the existing personalization methods aforementioned is likely to fail when training images do not represent the same instance, but rather encompass a distribution sharing certain, yet challenging-to-articulate, commonalities. Additionally, existing personalization methods often result in limited diversity and variation during generation (LABEL:fig:comparison). Since the associated token is fixed, these methods will typically learn a token that is either overfitted to a combination of visual features, or learn a token that is overly generalized, which introduces more randomness into the uncontrollable diffusion process, thereby failing to follow desired visual attributes in generated images.
In this work, we propose DreamDistribution, a prompt distribution learning approach on T2I diffusion model for various downstream tasks (LABEL:fig:teaser). Our proposed solution has three key components (LABEL:fig:method). First, to adapt a pretrained fixed T2I model, instead of fine-tuning diffusion model parameters, our method builds on prompt tuning (Zhou et al., 2022a; b), where we use soft learnable prompt embeddings with the flexibility to concatenate with text, to associate with the training image set. This design have several advantages: (1) It prevents catastrophic forgetting of the pretrained model, enabling it to learn an almost infinite variety of target prompt distributions using the same T2I diffusion model. (2) It is highly efficient in terms of parameters, requiring only the prompt itself as the learnable element. (3) The learned prompts remain within the semantic space of natural language, offering text-guided editing capabilities and generalizing to other pre-trained diffusion models, such as text-to-3D. (4) The learned distribution increased flexibility in managing variations. Second, we introduce a distribution of prompts to model various attributes described by reference images at a broader level. The prompt distribution is modeled by a set of learnable prompt embeddings to associate with the training image set as a whole. The learned prompt distribution can be treated as a distribution of learned “descriptions” of the reference images and should be able to model the commonalities and variations of visual attributes, e.g., foreground, style, background, texture, pose. During inference, we sample from the prompt distribution, which should have a similar semantic meaning, understood by the downstream denoising network, to produce in-distribution outputs with appropriate variations. Lastly, to effectively optimize the set of soft prompts that models the distribution, we apply a simple reparameterization trick (Kingma & Welling, 2013) and an orthogonal loss to update the prompts at token embedding space simultaneously and orthogonally.
We first demonstrate the effectiveness of our approach in customizing image generation tasks (LABEL:sec:experiments). By taking a small set of images of interest as training images, we demonstrate that our approach can generate diverse in-distribution images where baseline methods fail to generate desired output. The diversity and the quality of our synthetic images are verified via automatic and human evaluation (LABEL:sec:quantitative). We show that the learned distribution holds the capability of text-guided editing, as well as further controllability such as scaling the variance and composition of distributions (LABEL:sec:manipulation). Next we highlight that the learned prompt distribution can be easily applied to other text-guided generation tasks such as pretrained text-to-3D models (LABEL:sec:text23d). Lastly we show the effectiveness of our method on personalized distribution generation through classification task with synthetic training data as a proxy (LABEL:sec:imagenet). In summary, our contributions are:
  • We propose a distribution-based prompt tuning method for diverse personalized generation by learning prompt distribution using T2I diffusion model.
  • Using a public available pretrained T2I diffusion model, we experiment our approach on customization T2I generation tasks and show that our approach can capture visual attributes into prompt distribution and can generate diverse in-distribution images that follows text-guided editing.
  • Further experiments show that our method is flexible in terms of diversity or mixing and easy to be adapted to other text-guided generation tasks.
  • We further quantitatively demonstrate the effectiveness of our approach using synthetic image dataset generation tasks as a proxy and also through automatic evaluation metrics and human evaluation.
2 Acknowledgment
This work was supported by the National Science Foundation (award 2318101), C-BRIC (one of six centers in JUMP, a Semiconductor Research Corporation (SRC) program sponsored by DARPA), the Army Research Office (W911NF2020053), and Amazon ML Fellowship. The authors affirm that the views expressed herein are solely their own, and do not represent the views of the United States government or any agency thereof.
Appendix A More Implementation Details
In all experiments, we use Stable Diffusion 2.1, which is the latest version. We use the default parameters, including 7.5 guidance scale and 50 denoising steps. For all visual results, we generate images in $768\times 768$ resolution, and for synthetic dataset experiments, we generate $256\times 256$ images to save time and resources. We provide a pseudocode for learning a prompt distribution in Algorithm 1.
Algorithm 1 Training prompt distribution
1:Set of reference images =$\mathcal{I}=\{\mathbf{x}_{i}\}_{i=1}^{N}$
2:Set of learnable prompts $\mathcal{P}^{K}=\{\mathcal{P}_{i}\}_{i=1}^{K}$
3:Text encoder $\mathcal{E}$, noise predictor $\bm{\epsilon}_{\theta}$, hyperparameter $\lambda$
4:Random initialize all learnable embeddings $\mathbf{V}$ in $\mathcal{P}^{K}$
5:for image $\mathbf{x}\in\mathcal{I}$ do
6:     Sample time step $t$
7:     Sample noise $\bm{\epsilon}\sim\mathcal{N}(0,\bm{I})$
8:     $\mathbf{x}_{t}\leftarrow\mathbf{x}$ with noise added based on $\bm{\epsilon}$ and $t$
9:     Compute $\bm{\mu_{c}}=\frac{1}{K}\sum_{i=1}^{K}\mathcal{E}\left(\mathcal{P}_{i}\right)$
10:     Compute $\bm{\sigma_{c}^{2}}=\frac{1}{K}\sum_{i=1}^{K}\left(\mathcal{E}\left(\mathcal{P}_{i}\right)-\bm{\mu_{c}}\right)^{2}$
11:     for $s\in[S]$ do
12:         Sample $\bm{\omega}_{s}\sim\mathcal{N}(0,\mathbf{I})$
13:         $\mathcal{L}_{s}\left(\mathcal{P}^{K}\right)=\left\|\bm{\epsilon}-\bm{\epsilon}_{\theta}\left(\mathbf{x}_{t},\bm{\mu_{c}}+\bm{\omega}_{s}\bm{\sigma_{c}},t\right)\right\|^{2}_{2}$
14:     end for
15:     $\mathcal{L}\left(\mathcal{P}^{K}\right)=\frac{1}{S}\sum_{s=1}^{S}\mathcal{L}_{s}(\mathcal{P}^{K})$
16:     $\mathcal{L}_{\text{ortho}}=\frac{1}{K(K-1)}\sum_{i=1}^{K}\sum_{j=i+1}^{K}\left|\langle\mathcal{E}(\mathcal{P}_{i}),\mathcal{E}(\mathcal{P}_{j})\rangle\right|$
17:     $\mathcal{L}=\mathcal{L}\left(\mathcal{P}^{K}\right)+\lambda\mathcal{L}_{\text{ortho}}$
18:     Update learnable embeddings $\mathbf{V}$ in $\mathcal{P}^{K}$ based on $\mathcal{L}$
19:end for
Figure 1: In general, with more prompts, the performance increases in terms of both quality and diversity.
Figure 2: Choice of $\lambda$ value between $1\times 10^{-4}$ and $1\times 10^{-2}$ generally achieves good balance of quantitative metrics.
Figure 3: Similar to number of prompts, increasing number of prompt tokens also shows increasing image quality and diversity.
Appendix B Evaluation Set
We show samples of reference images from our evaluation set in Figure 4. Each row shows 8 samples reference images from the same set.
Figure 4: Samples of reference images from our evaluation set. Numbers on the right represent the number of images in each set.
Appendix C Additional Result
C.1 Diverse Image Instance Generation
We show additional image generation results of our method in Figure 5 using the reference images from our evaluation set. Each row is generated using reference images of the corresponding row in Figure 4. We also show additional examples in Figure 6 that demonstrate the ability of our method to capture visual similarities in reference images and generate diverse images, specifically on the attributes that are shown different in reference images.
Figure 5: Samples of generated image results using reference images from the evaluation set. Each row is generated using reference images of the corresponding row in Figure 4.
C.2 Text-guided Editing
We show more results on the ability of our method to generate diverse images with text-guided editing in Figure 7 using different sets of reference images.
Figure 6: More results that shows the ability of our method to capture same visual attributes and generate diverse images specifically on the attributes that are shown different in reference images. The top example shows reference images with same shape outline but different foreground texture and pattern, and our method also generates the same shape with diverse new foreground patterns. The bottom example shows reference images of same object in different background, and our method generates the same object in diverse background as well. We additionally show comparison with a naive adaptation of Textual Inversion, where we add a Gaussian noise with 0.5 variance to the learned tokens for generation, and the result shows that although it may increase diversity due to introduction of extra randomness, it fails to capture the common attributes that should not be varied in both examples, such as constant shape and background in first example and color and appearance in second example.
Figure 7: More results on text-editability of our methods. Left column shows samples of reference images to learn distribution $D^{*}$, right columns are generated results using prompts sampled from corresponding text-edited distribution.
C.3 Scaling Variance for Diversity Control
We show more results on generating images from prompt distribution with scaled standard deviations in Figure 8
Figure 8: More results on scaling standard deviation of learned prompt distribution.
C.4 Composition of Distribution
In Figure 9 we show more results on composition of two different learned prompt distributions with various weights.
Figure 9: More results on composition of multiple prompt distributions using different weights.
C.5 Same Instance Personalization
Our method solves a more generalized task compared to same instance personalization, where reference images is one or more images of the same instance and the generation is expected to follow user provided text prompts of different contexts. Although reconstruction of the same instance is not our main focus, we still show that our method can achieve on par results on this task compared with baselines, using DreamBooth dataset (Ruiz et al., 2022) with identical prompts and evaluation settings. The quantitative results are show in Table 1.
[…]
DreamBooth 0.80 0.31 0.67
Textual Inversion 0.78 0.26 0.57
Custom Diffusion 0.80 0.36 0.54
Ours 0.80 0.31 0.59
Table 1: Experiment results of personalized generation on DreamBooth dataset.
C.6 Naive Adaptation of Baselines
We compare our approach with a naive adaptation of baseline methods, where the training set is randomly divided into 4 subsets, and a generation model is trained on each subset. During inference, the final output is a mixture of results from the models trained on these different subsets. Based on the evaluation results shown in Section C.6, our method outperforms this naive adaptation of three baseline methods on all metrics. Moreover, this naive adaptation takes significantly more extra disk space since it requires the storage of multiple sets of model weights.
{NiceTabular}
[…]
DreamBooth 0.80 0.26 0.44 1.00 0.83 232.13
Textual Inversion 0.80 0.25 0.44 0.78 0.66 243.93
Custom Diffusion 0.75 0.27 0.39 0.62 0.57 268.48
Ours 0.84 0.29 0.50 1.59 0.93 215.15
Table 2: Comparison of our methods with naive adaptations of baseline methods.
C.7 Exploring with Different Granularity of Concepts
We experiment on image sets at different class granularity to show that our method is capable of capturing distribution at different levels of concepts. Specifically, we composed 3 different reference image sets, forced from ImageNet images: French Bulldog, dogs, and quadruped animals, where dogs set includes images from different ImageNet dog species such as Bulldog, Saluki, Chihuahua, etc., and animals set includes images from different ImageNet quadruped classes, such as tiger, cat, bison, fox, etc. We run DreamDistribution on these image sets respectively. The result visualization is shown in Figure 10. For French Bulldog, our method is able to generate different instances of French Bulldogs with different fur colors and appearances. For dogs reference set, our method is able to generate different dog species with mixed attributes. For quadrupeds reference set, our method can generate a variety of quadruped animals, including non-existing animals that show a mix of attributes from different animals. Quantitative analysis presented in Section C.7 demonstrates that at the most granular level, such as French Bulldog, our method significantly outperforms the personalization baselines. However, at coarser-grained levels, such as Dogs or Quadrupeds, the advantage of our method diminishes. This could be attributed to the increased diversity of the reference images, which makes it challenging for our method to effectively capture a coherent or meaningful distribution, leading to a higher likelihood of sampling outliers from the learned prompt distribution during generation.
Figure 10: Visualization of reference and generated images using different methods at different class granularity.
{NiceTabular}
[…]
Bulldog Textual Inversion 0.82 0.49 0.10 0.12 98.76
DreamBooth 0.85 0.56 0.02 0.04 100.24
Custom Diffusion 0.84 0.54 0.09 0.14 98.39
Ours 0.85 0.61 0.23 0.34 88.81
[…]
DreamBooth 0.76 0.20 0.68 0.44 211.34
Custom Diffusion 0.77 0.21 1.00 0.64 179.51
Ours 0.77 0.21 0.93 0.64 173.74
[…]
DreamBooth 0.69 0.16 2.35 0.23 228.39
Custom Diffusion 0.70 0.15 2.20 0.21 217.58
Ours 0.71 0.18 1.89 0.24 210.93
Table 3: Comparison of our method with baselines on feature-based metrics over different granularities of reference images.
C.8 Experiment with Small Reference Set
We conduct experiments on image sets with only four reference images to simulate scenarios where users are unable to provide a larger set for better distribution capture. Our method is compared against the personalization baselines used in previous sections. Visualizations of the generated results are shown in Figure 11, which shows that our method can still generate results with more diversity and at the same time within the distribution of reference images.
Figure 11: Visual results of training on only 4 reference images.
C.9 3D Generation Diversity
Our learned prompt distribution can extend to text-to-3D generation pipelines, such as MVDream Shi et al. (2023b), which utilize text-to-image models as their backbone, enabling more diverse 3D asset generation. As illustrated in Figure 12, unlike the Textual Inversion baseline, which relies on a fixed learned token to prompt the text-to-3D model, our method samples from the learned prompt distribution, resulting in greater diversity in the generated 3D assets.
Following Corso et al. (2023), we perform a quantitative analysis of in-batch cosine similarities among features of rendered images. Specifically, we compute the average pairwise cosine similarity of DINO and CLIP-I features within each image set. The results shown in Table 4 demonstrate that 3D assets generated using our learned prompt distribution exhibit significantly greater diversity compared to those generated using Textual Inversion’s learned token, which aligns with visual results. All samples are generated without using random seeds.
Figure 12: Qualitative comparison of diverse 3D generation.
[…]
Textual Inversion 0.63 0.31
Ours 0.53 0.14
Table 4: Quantitative comparison of 3D generation diversity using averaged pairwise cosine similarity of features of rendered images. Lower similarity means the generated set is more diverse.
Appendix D Ablation study
Number of Prompts We additionally ablate $K$, the number of prompts in personalized generation. We randomly select 4 sets of reference images from our evaluation set and compute the average performance based on automatic quality and diversity metrics introduced in main Section 4.1. In Figure 1, we show the effect of $K$ in terms of both generation quality and diversity. We observe a positive correlation between the performance (in terms of both quality and diversity) and the number of prompts. More prompts offer more flexibility for our methods to model a better distribution of prompts, thus enabling the model to encapsulate content better (quality) and adapt to various nuances of the training images (diversity).
Orthogonal Loss We also test different value of weight of orthogonal loss added during prompt distribution training. As shown in Figure 2, our choice of $\lambda$ (x-axis) can achieve good balance on quantitative metrics.
Number of Prompt Token Vectors We evaluate the impact of varying the number of token vectors $M$ used in each learnable prompt. As shown in Figure 3, similar to number of prompts, we observe a positive correlation between performance and the number of learnable prompt tokens. However, this effect is less pronounced compared to the impact of changing the number of prompts, and increasing the number of learnable tokens results in longer training times.
Appendix E Synthetic dataset
Figure 13: Column 1 & 2: Top-1 and top-5 accuracy on ImageNet validation set versus using different number of training images to train prompt distribution. Column 3 & 4; Top-1 and top-5 accuracy on ImageNet validation set versus percentage of synthetic images added to the real training set.
{NiceTabular}
[…]
Top1 Top5 Top1 Top5 Top1 Top5 Top1 Top5 Top1 Top5
10 49.8 73.8 48.5 71.4 27.7 48.8 57.1 78.1 15.3 44.4
100 56.7 79.8 62.0 81.5 30.4 50.9 54.7 78.4 18.8 55.0
500 64.3 84.0 61.7 81.6 25.2 45.8 53.0 74.8 15.7 50.4
full(1300) 63.6 82.9 60.5 80.0 18.6 36.3 43.5 67.8 11.6 45.0
Table 5: ImageNet classification accuracy on different real test sets by training a classifier on synthetic ImageNet (IN) generated by DreamDistribution using different number of real training images per class. Our experiments show that using 500 training images per class yields the highest final classification accuracy on real validation set.
E.1 More Analysis on Synthetic Dataset
Number of training images
We experiment with different number of training images. We use randomly selected 10, 100, 500 images per class, as well as all ImageNet training images to train our learnable prompts and generate same size synthetic dataset. From results shown in Appendix E and Figure 13 column 1 & 2, we found that using about 100-500 images per class (7%-38% of the real training set) would be enough to reach high classification accuracy on real validation set, while using more data would not further improve accuracy. For validation sets of ImageNet variants, less training data would obtain higher accuracy due to the domain gap between the real training set and different validation sets.
Mixing synthetic data with real training data
We also experiment with mixing different sizes of synthetic image data with the real training images. We mix additional 20%, 40%, 60%, 80% and 100% synthetic data with real training data, where 100% means the size of the mixed dataset is twice of the size of the ImageNet training set, and the ratio of the number of real images to the number of synthetic images is approximately 1:1. As shown in Figure 13 column 3 & 4, adding more synthetic data would improve the accuracy on ImageNet validation set. On the validation sets of different domains, however, adding more synthetic data would not show significant improving trends on accuracy.
E.2 Implementation detail
For training on ImageNet dataset, we use the same training hyperparameters except for reducing the number of learnable prompts to 10 per class. We train for 5 epochs for training prompt distribution and 300 epochs for training ResNet-50 using generated or mixed dataset. All results are averaged over 3 runs of training using the generated or mixed dataset. For ImageNet-R and ImageNet-A, we only evaluate on the overlapping classes with ImageNet-100.
E.3 More Visual result
We show some generated training images using our method and compare them with generated images using baseline methods in Figure 14, Figure 15, Figure 16, Figure 17, Figure 18, Figure 19. Compared to the images generated using baseline methods, non-learning methods could generate images with diversity, however, the appearances are generally far from real images. Learning based personalization baselines can generate images with high fidelity, but with very limited diversity. Our method in contrast can generated images that looks more like real image samples and at the same time with significant diversity.
Appendix F Limitations
Despite the ability of our method to generate diverse novel in-distribution images, it does have certain limitations. Specifically, our method may struggle to capture visual features when the number of training images is limited and very diverse. Moreover, the Gaussian distribution assumption could be overly restrictive depending on the training images and the text encoder’s latent space. In the future, we hope to find a more robust approach to learning distributions from a few, highly diverse images, with more accurate assumptions and resilient distribution forms.
Figure 14: Visualization of generated images on ImageNet ambulance.
Figure 15: Visualization of generated images on ImageNet cabbage.
Figure 16: Visualization of generated images on ImageNet honeycomb.
Figure 17: Visualization of generated images on ImageNet lorikeet.
Figure 18: Visualization of generated images on ImageNet papillon.
Figure 19: Visualization of generated images on ImageNet pirate ship.